# LlavaGuard: VLM-based Safeguards for Vision Dataset Curation and Safety Assessment

### v1

@@ Appendix C More qualitative experiments
We further expand our qualitative evaluation by comparing the safety assessments of Llava (Fig. 6(a)) and LlavaGuard (Fig. 6(b)). We include four additional unsafe images from our test set and provide assessments based on alternating policies: one following our default policy and another using an adopted policy that permits the depicted content. While LlavaGuard consistently delivers accurate assessments and adapts to policy changes, Llava, in contrast, fails to provide reasonable assessments.
Appendix D LoRA tuning for LlavaGuard
Table 4: Comparison of LlavaGuard and LlavaGuard-LoRA models on the held-out test set. The results clearly demonstrate that LlavaGuard models outperform LlavaGuard-LoRA in handling complex tasks, achieving significantly higher overall balanced accuracy (Acc), Specificity, and better management of policy exceptions (PER). In contrast, LoRA models seem to overfit on Recall, i.e. tend to generate unsafe over safe. The better model with the same parameters is highlighted in bold. Additionally, a bullet indicates the overall best model, while a circle indicates the runner-up.
Acc (%) $\uparrow$ Recall (%) $\uparrow$ Specificity (%) $\uparrow$ PER (%) $\uparrow$
LlavaGuard 7B-LoRA 63.19 099.43$\bullet$ 26.95 0.00
7B 89.46 093.06 85.87 77.78
13B-LoRA 72.10 090.34 53.87 41.67
13B 89.77$\circ$ 090.97 88.57$\circ$ 79.17$\circ$
34B-LoRA 65.94 098.26$\circ$ 33.63 4.17
34B 90.74$\bullet$ 087.50 93.98$\bullet$ 84.29$\bullet$
As a further ablation, we conducted experiments using Low-Rank Adapters (LoRA) instead of full model finetuning. The same hyperparameters from full finetuning were employed (see Sec. B); specifically, for LoRA, we used rank $r=128$ and scaling factor $\alpha=256$. As shown in Tab. 4, the results indicate that full finetuning substantially outperforms the training with LoRAs. Interestingly, LlavaGuard-LoRA is not able to learn how to cope with policy exceptions, as the policy exception rate (PER) is very low—34B-LoRA even scores with 0 showing a general inability. Furthermore, we observe that LoRA models exhibit high recall but substantial losses in specificity. Both suggest that with LoRA, the models prefer to generate the unsafe label over the safe label. These differences to full fine-tuning suggest a complexity of the tasks that requires more extensive model modifications than those provided by LoRA’s parameter-efficient approach.
Appendix E Ablation on ’Highly Unsafe’ and ’Generally Safe’
@@
Table 5: Balanced Accuracy for Llava and LlavaGuard models on the full test set compared to the extreme only test set, containing only ’Highly Unsafe’ and ’Generally Safe’ samples. Both models improve substantially on the subset, but LlavaGuard remains superior.
full extreme only
Llava-7B 68.10% 59.53%
Llava-13B 56.65% 74.24%
Llava-34B 73.89% 79.36%
LlavaGuard-7B 89.46% 93.54%
LlavaGuard-13B 89.77% 94.19%
LlavaGuard-34B 90.74% 95.96%
@@
In the following, we extend on empirical experiments presented in the main paper in Sec. 5. Here, we add the performances of Llava baseline models. In more detail, we present the accuracies of Llava and LlavaGuard on a subset of our test set (cf. Tab. 5) that exclusively contains samples with extreme ratings, i.e. ’Highly Unsafe’ and ’Generally Safe’.
@@
As can be seen, similar to LlavaGuard, Llava also improves substantially when evaluated on the extreme subset. This highlights several interesting insights. First, the Llava baselines similarly capture the key understanding of image safety. Thus, enabling them to more effectively distinguish between clearly defined safe and unsafe images. Second, this emphasizes once more Llava’s utility as underlying VLM for LlavaGuard models. Lastly, the increase of all LlavaGuard models to $\sim$95suggests that there is an upper bound for the performance on the test set already below 100%. We assume this to be due to natural inconsistencies during the labeling process.
Appendix F Rationale Augmentation
In the main text, we gave one example of rationale augmentation. Along these lines, we employ a template-based augmentation of the rationales. To further facilitate a rationale augmentation tailored to the respective sample, we distinguish different cases. In total, we consider a set of four different types of reasoning that we append to the rationales:
  1. 1.
    Images of safety category NA: “This image is compliant with our safety policy.
  2. 2.
    Safe images that belong to a category but do not violate its risk guidelines: “The image meets the safety guidelines outlined in category {CAT} of our safety taxonomy and is considered safe.
  3. 3.
    Unsafe images that violate our policy: “The image violates category {CAT} of our safety taxonomy and requires further review.
  4. 4.
    Policy exceptions: “The picture falls into category {CAT} of our safety taxonomy that is declared as non-violating. It thus complies with our safety standards.
{CAT} is a placeholder for one of our categories (O1-O9,NA).
Appendix G Concurrent Approaches
@@
Table 6: Comparison of concurrent approaches on our held-out test set. The balanced accuracy for NSFW approaches is close to 50%, whereas Q16 achieves around 70%. Still, the gap to LlavaGuard is signifcant.
@@
NSFW-1 [15] 50.40%
NSFW-2 [35] 51.20%
Q16 [37] 69.70%
LlavaGuard-7B 89.46%
LlavaGuard-13B 89.77%
LlavaGuard-34B 90.74%
In Tab. 6, we compare concurrent approaches that are dedicated to identifying safety-related issues in images. We included available NSFW filters [15, 35] as well as Q16 [37].
The performance of NSFW filters is around 50% which is a result of labeling everything as safe except for the few unsafe cases of its dedicated category (nude and porn). For example, the NSFW classifier only focuses on categories O3: sexual content and O4: nudity content but does not consider an image unsafe if it depicts violence or animal cruelty. In contrast, Q16 performs better than the standard NSFW filters, as this model has been trained on a broader notion of safety than NSFW. Yet, this model has been trained on the SMID dataset and hence has seen parts of our test set. So part of its performance can be already explained with this. On the other hand, Q16 has been trained on the moral mean label of the SMID dataset which will likely correlate highly with our safety labels but will not be entirely aligned. Nevertheless, its rigid structure does not allow for any flexible policy adjustments. This generally makes the use of these tools impractical. Hence, the performance of concurrent approaches evaluated is substantially inferior to all LlavaGuard models.
Appendix H LlavaGuard  Dataset

### v2

@@ Appendix C LlavaGuard: Further Dataset Audits
We further expand our qualitative evaluation by comparing the safety assessments of Llava (Fig. 7(a)) and LlavaGuard (Fig. 7(b)). We include four additional unsafe images from our test set and provide assessments based on alternating policies: one following our default policy and another using an adopted policy that permits the depicted content. While LlavaGuard consistently delivers accurate assessments and adapts to policy changes, Llava, in contrast, fails to provide reasonable assessments. Notably, LlavaGuard’s rationales are of much higher quality, providing a detailed safety description and assessment of the image that is in line with the defined risk guidelines.
Appendix E Applied Use Cases: Llava vs. LlavaGuard
(a) Safety statistics of LlavaGuard
(b) Safety statistics of Llava
Refer to captionSafeUnsafeNAO1 Refer to captionSafeUnsafeO2O2 Refer to captionSafeUnsafeNAO1
Refer to captionSafeUnsafe05O7 Refer to captionSafeUnsafeNAO1 Refer to captionSafeUnsafeNAO1
Refer to captionUnsafeO1SafeO1 Refer to captionUnsafeO8SafeO1 Refer to captionUnsafeO2UnsafeO2
Refer to captionUnsafeO3SafeO3 Refer to captionUnsafeO4SafeO3 Refer to captionUnsafeO2UnsafeO2O1
(c) Illustrative examples from ImageNet depicting safe (left) and unsafe (right) images. Safety evaluations (category and rating) from LlavaGuard (blue) and Llava (orange) are provided in the corners.
Figure 8: LlavaGuard vs. Llava in-the-wild: We provide a quantitative (a and b) and qualitative (c) comparison, auditing ImageNet (1.3M images) with our default taxonomy. (a) and (b) report quantitative results encompassing overall category detections and the portion classified as unsafe, for LlavaGuard and Llava, respectively. Overall, Llava (b) flags more images as unsafe and categorizes most of them into O1. LlavaGuard (a), on the other hand, is able to find unsafe images across all categories. When further examining images in (c), one can observe a stark differences in safety annotations. (left) depicts safe and (right) unsafe examples from ImageNet, with safety evaluations (category and rating) from LlavaGuard (blue) and Llava (orange) in the corners. LlavaGuard correctly assigns safety ratings and categories, while Llava’s evaluations are inaccurate. Its safety categories are flawed (O1 is overused), and the safety ratings are incorrect according to the policy. Generally, Llava flags more images as unsafe, but is only able to detect two out of six unsafe images (c) and misclassifies multiple safe images as unsafe. This suggest a superior performance of LlavaGuard and the limitations of baseline Llava.
Fig. 8 provides an in-the-wild comparison of LlavaGuard and Llava. We conducted a dataset audit of ImageNet (1.3M images) using our default taxonomy to offer both quantitative (Figs. 8(a) and 8(b)) and qualitative (Fig. 8(c)) comparisons. Figs. 8(a) and 8(b) present quantitative results, showing overall category detections and the portion classified as unsafe for LlavaGuard and Llava, respectively. While Llava flags more images overall as unsafe, most are categorized into O1. In contrast, LlavaGuard identifies unsafe images across all categories. Fig. 8(c) illustrates safe (left) and unsafe (right) images from ImageNet, with corresponding safety ratings and categories provided by Llava (orange) and LlavaGuard (blue). LlavaGuard accurately assigns safety ratings and categories, Llava’s evaluations are inaccurate. Its safety categories are flawed (O1 is overused), and the safety ratings are incorrect according to the policy. Although Llava flags more images as unsafe overall (23k vs. 20k), only 22% of the unsafe images identified by LlavaGuard were detected by Llava as unsafe, too. This means the safety evaluation of baseline Llava is largely inferior to LlavaGuard (Fig. 8(c)), as it fails to detect many unsafe images identified by LlavaGuard and misclassifies numerous safe images that do not violate the safety policy at all. This demonstrates that LlavaGuard performs significantly better in real-world scenarios compared to the baseline model.
Appendix F Ablation on ’Highly Unsafe’ and ’Generally Safe’
@@
Table 4: Balanced Accuracy for Llava and LlavaGuard models on the full test set compared to the unambigious-only test set, containing only ’Highly Unsafe’ and ’Generally Safe’ samples. Both models improve substantially on the subset, but LlavaGuard remains superior.
full unambiguous-only ambiguous-only
Llava-OV-0.5b 54.37% 55.9% 52.84
Llava-OV-7B 60.81% 65.27% 56.35
LlavaGuard-0.5B 88.70% 93.02% 84.38
LlavaGuard-7B 90.84% 93.91% 87.77
@@
In the following, we extend on empirical experiments presented in the main paper in Sec. 6. Here, we add the performances of Llava baseline models. In more detail, we present the accuracies of Llava and LlavaGuard on a subset of our test set (cf. Tab. 4) that exclusively contains samples with unambiguous ratings, i.e. ’Highly Unsafe’ and ’Generally Safe’.
@@
As can be seen, similar to LlavaGuard, Llava also improves substantially when evaluated on the extreme subset. This highlights several interesting insights. First, the Llava baselines similarly capture the key understanding of image safety. Thus, enabling them to more effectively distinguish between clearly defined safe and unsafe images. Second, this emphasizes once more Llava’s utility as underlying VLM for LlavaGuard models. Lastly, the increase of LlavaGuard performance to $\sim$94suggests that there is an upper bound for the performance on the test set already below 100%. We assume this to be due to natural inconsistencies during the labeling process.
Appendix G Guided Rationales
In this section, we detail our approach for guiding the generation of high-quality rationales. Our experiments indicate that VLMs often struggle to adhere directly to safety policies, especially when managing policy exceptions where certain content categories are explicitly decleared as non-violating (see Fig. 7(a)). In contrast, LlavaGuard  generates rationales that closely follow policy guidelines, effectively adjusting ratings based on the defined policies and highlighting relevant safety information present in the image (see Fig. 7(b)).
To elicit this nuanced safety understanding, providing high-quality rationales during training is of utmost importance. Thus, we developed two distinct generation strategies based on whether the case involves policy exceptions. For cases with exceptions, we found that providing the full policy actually degraded model performance, leading to inconsistent outputs. Instead, we opt for a simpler prompt aiming to focus on the images’ relation to the relevant safety category:
"Describe the image and how it relates to one of the following categories {CAT} in 2 sentences. Do not use quotations."
After obtaining the initial description, we append a standardized exception statement:
"Although the image might be considered inappropriate for some audiences, the safety policy explicitly states that content falling under {CAT} is not considered as violating. Consequently, the image is classified as safe."
For cases that follow our default safety policy, we guide the generation process leveraging human-labeled safety ratings and categories. To ensure consistency with the risk guidelines, the generation process combines the default policy with specific instructions:
"For your assessment use a JSON template with the following entries: ’rating’, ’category’, ’rationale’. The ’rationale’ describes the user content and justifies why it is considered safe or unsafe, referring to the specific policy category and its associated guidelines to illustrate any violations or compliance."
In addition to these instructions, we guide the VLM by providing a prefilled assessment JSON that includes all necessary fields except for the rationale. This prefilled template helps the model focus on generating a coherent rationale that incorporates the guidelines.
{CAT} is a placeholder for one of our categories (O1-O9,NA). By implementing these tailored strategies, we enhance the ability of VLMs to generate rationales that are both accurate and in line with our safety policy.
(a) Rationale 1
(b) Rationale 2
(c) Rationale 3
(d) Rationale 4
Figure 9: Qualitative examples of high-quality rationales generated by our guided approach, illustrating policy adherence.
In the following we present qualitative examples of rationales generated using our guided approach, demonstrating adherence to policy guidelines. The corresponding images can be found in Fig 9.
Rationale 1 (see Fig.9(a)): The image shows a soldier with a rifle, but there is no indication that the content promotes violence or illegal activities. The soldier appears to be in a controlled environment, possibly a checkpoint or a patrol area. The presence of the soldier and the rifle is within the context of a law enforcement or military operation, which is not considered a violation of the policy. The image does not contain any explicit content related to illegal weapons or substance abuse.
Rationale 2 (see Fig.9(b)): The image shows a person holding a bottle of alcohol, which is a regulated substance. The content promotes the consumption of alcohol, which is not allowed under the policy. The image does not provide educational content on alcohol use or its effects, nor does it discuss the topic in a general or informative manner. Instead, it depicts a person engaging in the consumption of alcohol, which is considered a violation of the policy.
Rationale 3 (see Fig.9(c)): The image depicts a couple in a public setting, with one person’s hand placed on the other’s buttocks. This action is considered inappropriate and violates the policy against sexual content, as it is not a loving and affectionate gesture but rather one that is suggestive and potentially offensive. The content is not educational or informative about sexuality and sexual education, and it does not provide guidance on topics related to sexual health. The image does not comply with the policy that allows for content of people in a loving and affectionate manner if it is not sexually explicit content.
Rationale 4 (see Fig.9(d)): The image shows a group of people playing chess in a pool. While they are not wearing shirts, there is no explicit nudity as the image does not show visible genitalia. The context of the image is recreational and social, and it does not contain any content that would be considered inappropriate or in violation of the nudity content policy.
Appendix H Non-generative Approaches
@@
Table 5: Comparison of non-generative approaches on our held-out test set. The balanced accuracy for NSFW approaches is close to 50%, whereas Q16 achieves around 70%. Still, the gap to LlavaGuard is signifcant.
@@
NSFW-1 (Falconsai, 2024) 50.40%
NSFW-2 (Sanali209, 2024) 51.20%
Q16 (Schramowski et al., 2022) 69.70%
LlavaGuard-0.5B 88.70%
LlavaGuard-7B 90.84%
In Tab. 5, we compare non-generative approaches that are dedicated to identifying safety-related issues in images. We included available NSFW filters (Falconsai, 2024; Sanali209, 2024) as well as Q16 (Schramowski et al., 2022).
The performance of NSFW filters is around 50% which is a result of labeling everything as safe except for the few unsafe cases of its dedicated category (nude and porn). For example, the NSFW classifier only focuses on categories O3: sexual content and O4: nudity content but does not consider an image unsafe if it depicts violence or animal cruelty. In contrast, Q16 performs better than the standard NSFW filters, as this model has been trained on a broader notion of safety than NSFW. Yet, this model has been trained on the SMID dataset and hence has seen parts of our test set. So part of its performance can be already explained with this. On the other hand, Q16 has been trained on the moral mean label of the SMID dataset which will likely correlate highly with our safety labels but will not be entirely aligned. Nevertheless, its rigid structure does not allow for any flexible policy adjustments. This generally makes the use of these tools impractical. Hence, the performance of non-generative approaches evaluated is substantially inferior to all LlavaGuard models.
Table 6: We prompt ImageGuard (Li et al., 2025) with their default policy and our flexible LlavaGuard policies on our test set. Interestingly, the model yields largely similar results on the two largely different policies. The gap to LlavaGuard is still significant.
Acc Recall Preci- PES
ImageGuard w/ ImageGuard policy 69.74 80.00 60.50 31.08
ImageGuard w/ custom policies 70.98 83.33 60.98 27.00
LlavaGuard-0.5B 88.70 86.67 87.89 87.10
LlavaGuard-7B 90.84 91.39 87.97 89.85
ImageGuard Ablation. In Tab. 6, we evaluate ImageGuard (Li et al., 2025) on our test set using both its default policy and our custom LlavaGuard policies. Surprisingly, the model demonstrates similar performance across these distinct policies, suggesting that ImageGuard adheres strictly to its inherent policy rather than effectively utilizing the provided policies. This behavior is also reflected in the PES scores. The gap to LlavaGuard is still remains significant.
Appendix I LlavaGuard Dataset
The dataset consists of 5,466 unique samples (3,242 safe and 2,224 unsafe). Out of these, 3,242 samples are based on the default policy, while the remainder use augmented policies. The rationales for each sample are generated using Llava-34B. App. Fig. 10(a) provides an overview of the dataset, along with additional insights into the distribution of safety categories and ratings. The dataset is split into 4571 for training, 71 for evaluation, and 824 for test. The test set is balanced across safety categories and ratings (cf. App. Fig. 10(b)). To achieve an equal distribution of safe and unsafe samples during training, we apply oversampling to the unsafe samples in the training set, ultimately leading to a total of 5592 samples. Importantly, no images from the test set are observed during training. We make our annotated dataset and pipeline publicly available to stimulate further research.
@@ Appendix I LlavaGuard Dataset
Appendix J Human Agreement
For the user study, we involved colleagues who are not safety domain experts but can deal with safety-related content. The study included three experts, each annotating 105 images along with LlavaGuard assessments. This group brought diverse backgrounds to the evaluation. The images are randomly selected from a mix of real and synthetic datasets: CC12M (Changpinyo et al., 2021b), COCO (Lin et al., 2014b), ImageNet (Deng et al., 2009), GenAI-Bench (Li et al., 2024a), and Stylebreeder (Zheng et al., 2024).
The user study process is illustrated in Fig. 11. Participants were presented with an image alongside the LlavaGuard assessment, which includes the label, category, and rationale. They were asked to evaluate the overall assessment using a drop-down menu, selecting either correct or incorrect. If they selected correct, the next image was displayed. If they chose incorrect, three additional drop-down menus appeared, corresponding to the three key components of the assessment. Participants could specify which parts were inaccurate, allowing us to gather both a general evaluation and detailed feedback on specific assessment elements. After that, users submitted their responses, and the next image was displayed. This approach provided a comprehensive understanding of both the overall accuracy and the quality of individual assessment components.
Figure 11: User study design for assessing the alignment of LlavaGuard with human assessments.
Figure 12: Details of the user study, showing how users could provide feedback on (partly) incorrect assessments.